OpenStudy (anonymous):

Let T be a function from R2 to R2 such that T(2,1)=(1,0) and T(1,-1)=(2,1). Find T(3,-4)

OpenStudy (anonymous):

I don't think you have enough information to predict where that would go. Any point in R2 would do, if you have no more specifications than you've given here.

OpenStudy (anonymous):

would the fact that T is a linear function help? because besides that, this is all that is given

myininaya (myininaya):

a(2,1)+b(2,-1)=(3,-4) We need to find constants a and b 2a+2b=3 => 2a+2b=3 1a-1b=-4 => -2a+2b=8 --------------- 0+4b=11 => b=11/4 So what is a?

OpenStudy (anonymous):

a=-5/4, but how did you come up with the equation a(2,1)+b(2,-1)

myininaya (myininaya):

oops that second pai rwas suppose to be (1,-1) my bad

myininaya (myininaya):

\[a(2,1)+b(1,-1)=(3,-4)\]

myininaya (myininaya):

my eyes were playing tricks on me

OpenStudy (anonymous):

ok i see now

myininaya (myininaya):

So we need to solve this for a and b

myininaya (myininaya):

one sec i'm figuring this as we go to i know that is the first step

myininaya (myininaya):

oh wait i think i'm fixing to work out the other details

OpenStudy (anonymous):

\[(3,-4)=a(2,1)+b(1,-1)\] 2a+b=3 a-b=-4 i get a=7/3 and b=19/3

myininaya (myininaya):

@zarkon I forgot what to do here after this part

myininaya (myininaya):

i know it is something stupid i'm forgetting

OpenStudy (anonymous):

not a prob, i actually dnt know what to do either

OpenStudy (zarkon):

T is linear

OpenStudy (zarkon):

I assume

OpenStudy (anonymous):

yup

OpenStudy (zarkon):

then T(av+bw)=aT(v)+bT(w)

OpenStudy (zarkon):

lost my connection to the site for a sec

myininaya (myininaya):

wait i got it i think don't say it zarkon please

myininaya (myininaya):

will you check me though lol

myininaya (myininaya):

\[T(3,-4)=T(a(2,1)+b(1,-1))=T(a(2,1))+T(b(1,-1))=aT(2,1)+bT(1,-1)\] \[=a(1,0)+b(2,1)\] Replace a with what you found Replace b with what you found The answer will be this with your replacements =(a+2b,b)

myininaya (myininaya):

Right Zarkon? :)

OpenStudy (zarkon):

yes

myininaya (myininaya):

:) I love you zarkon your the best platonic love of course

OpenStudy (anonymous):

that's what i got too, but the teacher marked it wrong was confused. and then T(2,1)=(1,0). with the (a+2b, b), if you let a=2 and b=1. T(2,1)=(4,1) which does not equal with what was given? why is it?

myininaya (myininaya):

so you used the a and b you found above?

myininaya (myininaya):

I didn't check if your arithmetic was right. I could.

myininaya (myininaya):

yeah i think i see something wrong

myininaya (myininaya):

2a+b=3 a-b=-4 ------- 3a=-1 a=-1/3

myininaya (myininaya):

-1/3-b=-4 -b=-4+1/3 -b=-12/3+1/3 -b=-11/3 b=11/3

OpenStudy (zarkon):

\[T(x,y)=\left(x-y,\frac{x-2y}{3}\right)\] if you want to find the image of any vector \(v=(x,y)\)

OpenStudy (anonymous):

here is what i did. given T(2,1)=(1,0) and T(1,-1)=(2,1) i let there be a X=(a,b) which is in R2 X=a(2,1)+b(1,-1) TX=aT(2,1)+bT(1,-1)=a(1,0)+b(2,1) so T(a,b)= (a+2b,b) T(3,-4)=(-5,-4) sorry I took so long,

OpenStudy (zarkon):

T(3,-4)=(7,11/3)

OpenStudy (zarkon):

use the a,b myininaya found

OpenStudy (anonymous):

i must have done some error in the arithmetic one the way

OpenStudy (zarkon):

or use my general formula ;)